# Graph sensitivity under join and decomposition

### v2

@@ 3.3. Joining an empty graph to a regular nonempty graph
Let $G$ be a regular graph with degree $d$ on $n$ vertices. Then $\overline{K}_{m}\vee G$ with $m\geq 2n-d-1$ satisfies $\sigma(\overline{K}_{m}\vee G)=m-n+d+1$.
@@
Let $k$ be the number of vertices of $G$ not included in $H$. Let $v\in H$. If $v\in\overline{K}_{m}$, then $\deg_{H}(v)=n-k$. We will show if $v\in G$, then $\deg_{H}(v)\geq m-n+d+1$. Since we assumed $m\geq 2n-d-1$, then $m-n+d+1\geq n>n-k$ and hence it will follow that $\Delta(H)\geq m+1-n+d$.
@@
When $G$ is a regular graph of degree $d$ with $n$ vertices and $m\geq 2n-d-1$, Theorem 3.18 yields $\sigma(\overline{K}_{m}\vee G)=m-n+d+1\geq n$ and hence $\{\overline{K}_{m}\vee G\}_{m=1}^{\infty}$ is sensitive, which is consistent with Corollary 3.8 but provides the exact sensitivity in exchange for the conditions on $G$ and $m$.
Example 3.22.
When $H$ is a regular graph of degree $d$ and order $\ell$ and $m\geq 2n\ell-d-1$, Theorem 3.18 yields that $\sigma(\overline{K}_{m}\vee nH)=m-n\ell+d+1$ for the family of generalized windmill graphs of type II. This implies $\{\overline{K}_{m}\vee nH\}_{m=1}^{\infty}$ is sensitive, as is consistent with Example 3.14.
@@ 4. Sensitivity and stable blocks
In this section we consider when a graph $G$ has a decomposition into sets called stable blocks. These are defined by Larson in [Larson-12] and allow isolating maximum independent sets and describing sensitivity.
@@
Note that if $S$ is a stable block, then $\mathrm{int}(S)$ is nonempty. Stable blocks were used to efficiently reduce independence number calculations to subgraphs [Larson-11]. We use them to bound sensitivity in Section 4.1, and in Theorem 4.19 to compute sensitivity precisely in terms of the blocks when there is a full partition of $V_{G}$ into such sets. We also define vertex identification and use it as a means to construct a stable block in Corollary 4.6, repeat this to bound the sensitivity of a rooted product in Section 4.2, and indicate how to construct a graph with a partition into stable blocks in Section 4.3. As an application we compute the sensitivity of a corona of two graphs and of families of successive corona graphs.
@@ 4.1. Single stable block
If $G[\mathrm{int}(H)]$ is empty then by definition $\sigma(G[\mathrm{int}(H)])$ is infinite so the conclusion of Proposition 4.2 still holds, but the statement is most useful when $G[\mathrm{int}(H)]$ is nonempty.
@@
For the star graph $S_{n}=K_{1}\vee\overline{K}_{n}$ with $\{v\}$ the distinguished vertex in $K_{1}$, and the cycle graph $C_{m}$ with $m\geq 3$ and any distinguished vertex $\{u\}$, the cycle-star graph [Clark-85] is $CS_{m,n}=(C_{m}+S_{n})/\{u,v\}$. It has $\alpha(CS_{m,n})=\left\lfloor\frac{m}{2}\right\rfloor+n$ and $\Delta(CS_{m,n})=n+2$. Since there is a maximum independent subset in $C_{m}$ not containing $u$ and $CS_{m,n}[\mathrm{int}(C_{m})]=C_{m}\setminus\{u\}$ is a nonempty path, we conclude that $\sigma(CS_{m,n})=\sigma(CS_{m,n}[\mathrm{int}(C_{m})])=1$ for $m\geq 5$ (so that the path has length at least $4$) and the family $\{CS_{m,n}\}_{n=1}^{\infty}$ is insensitive.
@@
The following definition used in [Larson-11] was originally given in [Zhang-90].
@@
It is possible that the empty set is the only maximum critical independent set, as is the case when $G=K_{n}$ for all $n\geq 3$. But if not, the Independence Decomposition Theorem of [Larson-11] produces a stable block.
@@
At the start of the proof of Theorem $2.4$ in [Larson-11], Larson proves that $I_{c}$ is a maximum independent subset of $G[X]$, so $\alpha(G[X])=|I_{c}|$.
@@
Larson also proves that the set $X$ is independent of the choice of maximum critical independent set [Larson-11] and that a maximum critical independent set can be found in polynomial time [Larson-07]. Hence when there exists a nonempty maximum critical independent set Proposition 4.11 provides an effective way to produce a stable block in a graph and use Proposition 4.2 to obtain an upper bound on sensitivity, which is most useful for proving a family is insensitive.
@@ 4.2. Rooted products and coronas
Given a graph $G$ of order $n$ and a set of graphs $\mathcal{H}=\{H_{1},\dots,H_{n}\}$ rooted at $v_{1},\dots,v_{n}$ respectively, each root can be identified with a vertex of $G$ to obtain the rooted product $G(H_{1},\dots,H_{n})$ of $\mathcal{H}$ by $G$. This construction was originally introduced in [Godsil-McKay-78] where the characteristic polynomial of its adjacency matrix was determined in terms of the characteristic polynomials for $H_{1},\dots,H_{n}$.
@@
For any graph $G$ with one or more distinguished vertices, using any $1$-cone $H$ whose interior is nonempty results in a wide variety of examples by applying Corollary 4.14 at the chosen vertices of $G$. For instance, if $H$ is the join of $K_{1}$ and $K_{n}$, $P_{n}$, or $C_{n}$ for $n\geq 2$, $n\geq 4$, or $n\geq 5$ respectively, as in Example 2.3, then the resulting graph will have sensitivity $1$, as will any tree built by successive vertex identification involving at least one path of length at least $4$. In particular the royal petunia graphs of  [Latyshev-Kokhas-23] contain a copy of $K_{1}\vee P_{n}$ as a stable block, and thus have sensitivity $1$.
If we identify a copy of $K_{1}\vee H$ rooted at the vertex in $K_{1}$ to each vertex in $G$ then the resulting rooted product is called the corona $G\odot H$ of $G$ and $H$. The corona operation was introduced in [Frucht-Harary-70] to produce a graph whose automorphism group could be easily described as a wreath product involving the automorphism groups of the component graphs.
@@
If $H$ is nonempty then $G\odot H=G(K_{1}\vee H)$ will satisfy the hypotheses of Corollary 4.14 and each copy of $K_{1}\vee H$ will be a stable block in $G\odot H$, so $\sigma(G\odot H)\leq\sigma(K_{1}\vee H)=\sigma(H)$. For example, the Type III windmill graphs in [Kooij-19] are $K_{m}\odot K_{n}$ and hence by Corollary 4.14 have sensitivity $\sigma(K_{m}\odot K_{n})=1$. Since for $m\geq 2$ we have
@@ 4.3. Stable block decomposition
In contrast to modules and a maximal modular partition, stable blocks do not have to exist and can intersect, so a partition of a given graph into stable blocks might not exist. However in the case where there is a full decomposition of a graph into stable blocks we obtain precise information about its sensitivity in terms of that of the blocks. We also describe an opposite construction that produces families to which Theorem 4.19 does apply. The main idea is that if the pieces do not interact too much then they determine the sensitivity.
@@
If the vertex set of a graph $G$ can be partitioned into stable blocks $V_{1},\dots,V_{n}$ then
@@
Let $V_{1},\dots,V_{n}$ be a partition of the vertex set of $G$ into stable blocks and for $1\leq i\leq n$ let $G_{i}=G[V_{i}]$ and let $A_{i}$ be a maximum independent set in $\mathrm{int}(G_{i})$. Suppose without loss of generality that $\sigma(G_{1})=\min\limits_{1\leq i\leq n}\sigma(G_{i})$. Clearly $\bigcup\limits_{i=1}^{n}A_{i}$ is an independent subset of $G$ of order $\sum\limits_{i=1}^{n}\alpha(G_{i})$, but by the pigeonhole principle any larger subset intersects some $G_{i}$ in at least $\alpha(G_{i})+1$ vertices, and thus is not independent in $G$. Hence $\alpha(G)=\sum\limits_{i=1}^{n}\alpha(G_{i})$.
@@
There is a natural way to construct graphs that decompose into stable blocks as in Theorem 4.19. Let $G$ be a graph, and for each vertex $v_{i}$ with $1\leq i\leq k$ of $G$ let $H_{i}$ be a graph corresponding to $v_{i}$ with vertex set $V_{i}$ and edge set $E_{i}$. Fix a maximum independent subset $A_{i}$ of each $H_{i}$. Now build a graph from $G$ and the $H_{i}$ so that the following properties hold:
@@
Observe that even when $H$ is empty, each copy of $K_{1}\vee H$ is a stable block in the corona $G\odot H=G(K_{1}\vee H)$. Since the copies of $K_{1}\vee H$ form a partition of the vertices of $G\odot H$, by Theorem 4.19 we have $\sigma(G\odot H)=\sigma(K_{1}\vee H)$, which equals $\sigma(H)$ when $H$ is nonempty. When $H$ is empty then $\sigma(K_{1}\vee H)=|V_{H}|$ by Example 2.4. ∎
@@ 5. Concluding remarks
In a generalized join $G[H_{1},\dots,H_{n}]$, each subgraph induced by $H_{i}$, $1\leq i\leq n$ is a module, i.e., a set of vertices that have the same neighbors outside $H_{i}$. Thus $H_{1},\dots,H_{n}$ forms a modular partition of the full graph, and conversely any modular partition of a graph can be used to express it as a generalized join. In general it is possible to partition the vertices of a graph into modules in multiple ways. The modular decomposition tree of a graph $G$ exhibits the nesting of its strong modules (see [HMSZ-24]), uniquely determines $G$ (see [Brandstadt-Le-Spinrad-99]), and the maximal proper modules in that tree form a partition of the vertex set of $G$ that is referred to as a maximal modular partition of $G$ (see [Habib-Paul-10]). See [Brandstadt-Le-Spinrad-99] for definitions, a list of several other terms that have been used for similar ideas in a variety of contexts, and precise statements of a modular decomposition theorem and related results. For a brief primer and an interesting new generalization see [HMSZ-24] and for further details and a reference list see [Habib-Paul-10].
When $G$ has more than one connected component $H_{1},\dots,H_{n}$ the unique maximal modular partition is the parallel decomposition $G=\overline{K}_{n}[H_{1},\dots,H_{n}]$, for which we described the sensitivity in Example 3.14. When the complement $\overline{G}$ has more than one connected component $H_{1},\dots,H_{n}$ then $G=K_{n}[H_{1},\dots,H_{n}]$ is the series decomposition, for which we described sensitivity in the lexicographic product case, $K_{n}\circ H$, in Example 3.15. When $G$ and $\overline{G}$ are connected, then for a maximal modular partition $H_{1},\dots,H_{n}$ of $G$ we can replace each $H_{i}$ with a single vertex and obtain a quotient $Q$ that has no nontrivial modules and allows us to write $G$ as a generalized join $G=Q[H_{1},\dots,H_{n}]$. For example, in the parallel and series decompositions the graphs $\overline{K}_{n}$ and $K_{n}$ are quotients of $G$ by $H_{1},\dots,H_{n}$. This suggests limitations on finding the sensitivity of a generalized join explicitly in terms of its components, since in full generality this would amount to finding the sensitivity of a general graph.
@@ 6. Acknowledgements
We thank Davis Bolt and Gabriel Ory for computing support carried out in SageMath [sagemath]. The Idaho State University Career Path Internship Program provided funding for that computing support and some work by Jacob Tolman.

### v3

@@ 3.3. Joining an empty graph to a regular nonempty graph
Let $G$ be a regular graph with degree $d$ on $n$ vertices. Then provided $m\geq 2n-d-1$ we have $\sigma(\overline{K}_{m}\vee G)=m-n+d+1$.
@@
Let $k$ be the number of vertices of $G$ not included in $H$ and let $v\in H$. If $v\in\overline{K}_{m}$, then $\deg_{H}(v)=n-k$. We will show if $v\in G$, then $\deg_{H}(v)\geq m-n+d+1$. Since we assumed $m\geq 2n-d-1$, then $m-n+d+1\geq n>n-k$ and hence it will follow that $\Delta(H)\geq m+1-n+d$.
@@
When $G$ is a regular graph of degree $d$ with $n$ vertices and $m\geq 2n-d-1$, Theorem 3.18 yields $\sigma(\overline{K}_{m}\vee G)=m-n+d+1\geq n$ and hence $\{\overline{K}_{m}\vee G\}_{m=1}^{\infty}$ is sensitive, which is consistent with Corollary 3.8 but provides the exact sensitivity in exchange for the conditions on $G$ and $m$. In particular, if $G=nH$ where $H$ is a regular graph of degree $d$ and order $\ell$, then $\sigma(\overline{K}_{m}\vee nH)=m-n\ell+d+1$ for the family of generalized windmill graphs of type II and $\{\overline{K}_{m}\vee nH\}_{m=1}^{\infty}$ is sensitive.
@@ 4. Sensitivity and stable blocks
In this section we consider when a graph $G$ has a decomposition into sets called stable blocks. These are defined by Larson in [21] and allow isolating maximum independent sets and describing sensitivity.
@@
Note that if $S$ is a stable block, then $\mathrm{int}(S)$ is not the empty set. Stable blocks were used to efficiently reduce independence number calculations to subgraphs in [20]. We use them to bound sensitivity in Section 4.1, and in Theorem 4.19 to compute sensitivity precisely in terms of the blocks when there is a full partition of $V_{G}$ into such sets. We also define vertex identification and use it as a means to construct a stable block in Corollary 4.6, repeat this to bound the sensitivity of a rooted product in Section 4.2, and indicate how to construct a graph with a partition into stable blocks in Section 4.3. As an application we compute the sensitivity of a corona of two graphs and of families of successive corona graphs.
@@ 4.1. Single stable block
The case of a star graph $G=K_{1}\vee\overline{K}_{n}$ with $H$ a proper subgraph containing the center vertex shows that the conclusion of Proposition 4.2 need not hold when $G[\mathrm{int}(H)]$ is empty.
@@
For the star graph $S_{n}=K_{1}\vee\overline{K}_{n}$ with $\{v\}$ the distinguished vertex in $K_{1}$, and the cycle graph $C_{m}$ with $m\geq 3$ and any distinguished vertex $\{u\}$, the cycle-star graph [7] is $CS_{m,n}=(C_{m}+S_{n})/\{u,v\}$. It has $\alpha(CS_{m,n})=\left\lfloor\frac{m}{2}\right\rfloor+n$ and $\Delta(CS_{m,n})=n+2$. Since there is a maximum independent subset in $C_{m}$ not containing $u$ and $CS_{m,n}[\mathrm{int}(C_{m})]=C_{m}\setminus\{u\}$ is a nonempty path, we conclude that $\sigma(CS_{m,n})=\sigma(CS_{m,n}[\mathrm{int}(C_{m})])=1$ for $m\geq 5$ (so that the path has length at least $4$) and the family $\{CS_{m,n}\}_{n=1}^{\infty}$ is insensitive.
@@
The following definition used in [20] was originally given in [31].
@@
It is possible that the empty set is the only maximum critical independent set, as is the case when $G=K_{n}$ for all $n\geq 3$. But if not, the Independence Decomposition Theorem of [20] produces a stable block.
@@
At the start of the proof of Theorem $2.4$ in [20], Larson proves that $I_{c}$ is a maximum independent subset of $G[X]$, so $\alpha(G[X])=|I_{c}|$.
@@
Larson also proves that the set $X$ is independent of the choice of maximum critical independent set [20] and that a maximum critical independent set can be found in polynomial time [19]. Hence when there exists a nonempty maximum critical independent set Proposition 4.11 provides an effective way to produce a stable block in a graph and use Proposition 4.2 to obtain an upper bound on sensitivity, which is most useful for proving a family is insensitive.
@@ 4.2. Rooted products and coronas
Given a graph $G$ of order $n$ and a set of graphs $\mathcal{H}=\{H_{1},\dots,H_{n}\}$ rooted at $v_{1},\dots,v_{n}$ respectively, each root can be identified with a vertex of $G$ to obtain the rooted product $G(H_{1},\dots,H_{n})$ of $\mathcal{H}$ by $G$. This construction was originally introduced in [12] where the characteristic polynomial of its adjacency matrix was determined in terms of the characteristic polynomials for $H_{1},\dots,H_{n}$.
@@
For any graph $G$ with one or more distinguished vertices, using any $1$-cone $H$ whose interior is nonempty results in a wide variety of examples by applying Corollary 4.14 at the chosen vertices of $G$. For instance, if $H$ is the join of $K_{1}$ and $K_{n}$, $P_{n}$, or $C_{n}$ for $n\geq 2$, $n\geq 4$, or $n\geq 5$ respectively, as in Example 2.3, then the resulting graph will have sensitivity $1$, as will any tree built by successive vertex identification involving at least one path of length at least $4$. In particular the royal petunia graphs of  [22] contain a copy of $K_{1}\vee P_{n}$ as a stable block, and thus have sensitivity $1$.
If we identify a copy of $K_{1}\vee H$ rooted at the vertex in $K_{1}$ to each vertex in $G$ then the resulting rooted product is called the corona $G\odot H$ of $G$ and $H$. The corona operation was introduced in [10] to produce a graph whose automorphism group could be easily described as a wreath product involving the automorphism groups of the component graphs.
@@
If $H$ is nonempty then $G\odot H=G(K_{1}\vee H)$ will satisfy the hypotheses of Corollary 4.14 and each copy of $K_{1}\vee H$ will be a stable block in $G\odot H$, so $\sigma(G\odot H)\leq\sigma(K_{1}\vee H)=\sigma(H)$. For example, the Type III windmill graphs in [18] are $K_{m}\odot K_{n}$ and hence by Corollary 4.14 have sensitivity $\sigma(K_{m}\odot K_{n})=1$. Since for $m\geq 2$ we have
@@ 4.3. Stable block decomposition
In contrast to modules and a maximal modular partition, stable blocks do not have to exist and can intersect, so a partition of a given graph into stable blocks might not exist. However in the case where there is a full decomposition of a graph into stable blocks we obtain precise information about its sensitivity in terms of that of the blocks. We also describe a construction that produces families to which Theorem 4.19 does apply. The main idea is that if the pieces do not interact too much then they determine the sensitivity.
@@
If the vertex set of a graph $G$ can be partitioned into stable blocks $V_{1},\dots,V_{n}$ such that $G[V_{i}]$ is nonempty for $1\leq i\leq n$, then
@@
Let $V_{1},\dots,V_{n}$ be a partition of the vertex set of $G$ into stable blocks such that for $1\leq i\leq n$ we have $G_{i}:=G[V_{i}]$ nonempty, and let $A_{i}$ be a maximum size independent set in $\mathrm{int}(G_{i})$. Suppose without loss of generality that $\sigma(G_{1})=\min\limits_{1\leq i\leq n}\sigma(G_{i})$. Clearly $\bigcup\limits_{i=1}^{n}A_{i}$ is an independent subset of $G$ of order $\sum\limits_{i=1}^{n}\alpha(G_{i})$, but by the pigeonhole principle any larger subset intersects some $G_{i}$ in at least $\alpha(G_{i})+1$ vertices, and thus is not independent in $G$. Hence $\alpha(G)=\sum\limits_{i=1}^{n}\alpha(G_{i})$.
@@
There is a natural way to construct graphs that decompose into stable blocks as in Theorem 4.19. Let $G$ be a graph, and for each vertex $v_{i}$ with $1\leq i\leq k$ of $G$ let $H_{i}$ be a nonempty graph corresponding to $v_{i}$ with vertex set $V_{i}$ and edge set $E_{i}$. Fix a maximum independent subset $A_{i}$ of each $H_{i}$. Now build a graph from $G$ and the $H_{i}$ so that the following properties hold:
@@
Observe that even when $H$ is empty, each copy of $K_{1}\vee H$ is a stable block that induces a nonempty graph in the corona $G\odot H=G(K_{1}\vee H)$. Since the copies of $K_{1}\vee H$ form a partition of the vertices of $G\odot H$, by Theorem 4.19 we have $\sigma(G\odot H)=\sigma(K_{1}\vee H)$, which equals $\sigma(H)$. ∎
@@ 5. Concluding remarks
In a generalized join $G[H_{1},\dots,H_{n}]$, each subgraph induced by $H_{i}$, $1\leq i\leq n$ is a module, i.e., a set of vertices that all have the same neighbors outside $H_{i}$. Thus $H_{1},\dots,H_{n}$ forms a modular partition of the full graph, and conversely any modular partition of a graph can be used to express it as a generalized join. In general it is possible to partition the vertices of a graph into modules in multiple ways. The modular decomposition tree of a graph $G$ exhibits the nesting of its strong modules (see [14]), uniquely determines $G$ (see [3]), and the maximal proper modules in that tree form a partition of the vertex set of $G$ that is referred to as a maximal modular partition of $G$ (see [15]). Reference [3] contains definitions, a list of several other terms that have been used for similar ideas in a variety of contexts, and precise statements of a modular decomposition theorem and related results. For a brief primer and an interesting new generalization see [14] and for further details and a reference list see [15].
When $G$ has more than one connected component $H_{1},\dots,H_{n}$ the unique maximal modular partition is the parallel decomposition $G=\overline{K}_{n}[H_{1},\dots,H_{n}]$, for which we described the sensitivity in Example 3.14. When the complement $\overline{G}$ has more than one connected component $H_{1},\dots,H_{n}$ then $G=K_{n}[H_{1},\dots,H_{n}]$ is the series decomposition, for which we described sensitivity in the lexicographic product case, $K_{n}\circ H$, in Example 3.15. When $G$ and $\overline{G}$ are connected, then for a maximal modular partition $H_{1},\dots,H_{n}$ of $G$ we can replace each $H_{i}$ with a single vertex and obtain a quotient $Q$ that has no nontrivial modules and can be used to write $G$ as a generalized join $G=Q[H_{1},\dots,H_{n}]$. For example, in the parallel and series decompositions the graphs $\overline{K}_{n}$ and $K_{n}$ are quotients of $G$ by $H_{1},\dots,H_{n}$. This suggests limitations on finding the sensitivity of a generalized join explicitly in terms of its components, since in full generality this would amount to finding the sensitivity of a general graph.
@@ 6. Acknowledgements
We thank Davis Bolt and Gabriel Ory for computing support carried out in SageMath [28]. The Idaho State University Career Path Internship Program provided funding for that computing support and some work by Jacob Tolman.